A bag of taffy contains three flavors: chocolate, maple, and cherry. There is an equal number of each flavour. What is the chance of reaching into the bag and getting a cherry-flavored piece? Show your answer as a fraction.
The chance is 1 out of 3. The fraction is ⅓ or 1 over three.
It is tiring to type this again but... hope this helps.
The chance of reaching into the bag and getting a cherry-flavored piece is 1 out of 3. The fraction is 1/3.
The probability is defined as the possibility of an event is equal to the ratio of the number of favorable outcomes and the total number of outcomes.
A bag of taffy contains three flavors: chocolate, maple, and cherry.
There is an equal number of each flavor
To determine the chance of reaching into the bag and getting a cherry-flavored piece
The probability that a the chance of reaching into the bag and getting a cherry-flavored piece = P(E)
There is an equal number of each flavor
Hence required probability P(E) = 1/3 = 0.33
Hence, the chance of reaching into the bag and getting a cherry-flavored piece is 1 out of 3. The fraction is 1/3..
Learn more about probability here:
brainly.com/question/11234923
#SPJ2